OpenStudy (anonymous):

Determine the value of n in the expansion of (a+b)^n in descending powers of a, it is found that two adjacent coefficients are each equal to 126.

OpenStudy (anonymous):

General term is \[nCr(a)^{n-r}(b)^{r}\]

OpenStudy (aum):

Construct a Pascal's triangle.

OpenStudy (anonymous):

I let term 1= the first 126 r+1=1 r=0 and then plugged r=0 into the r in the general term\[126=nC0(a)^{n-0}(b)^{0}\] when simplified i get 126=a^n

OpenStudy (anonymous):

yea i did that i got n=9, but i wanted know if there was a faster way

OpenStudy (aum):

\[ ^nC_r = ^nC_{r+1} = 126 \]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[\frac{ n! }{ (n-r)!r! }=\frac{ n! }{ (n-r+1)!(r+1)! }=126\]

OpenStudy (anonymous):

now how do i go about solving this abomination?

OpenStudy (aum):

Divide the first by the second and equate it to 1. Find r in terms of n. Then put r in the first one equate it to 126 and solve for n.

OpenStudy (anonymous):

What do you mean by the first and the second? first=nCr? second= nCr+1?

OpenStudy (aum):

yes.

OpenStudy (aum):

\[ \frac{ n! }{ (n-r)!r! } \div \frac{ n! }{ (n-r+1)!(r+1)! }=1 \]

OpenStudy (anonymous):

okay

OpenStudy (aum):

\[ \frac{ n! }{ (n-r)!r! } \div \frac{ n! }{ (n-r+1)!(r+1)! }=1 \\ \frac{ \cancel{n!} }{ (n-r)!r! } \times \frac{ (n-r+1)!(r+1)! }{\cancel{n!}} = 1 \\ \]

OpenStudy (anonymous):

yea that what i got, now i dont know what to do

OpenStudy (anonymous):

\[\frac{ n-r+1)!(r+1)! }{ (n-r)!r!}=1\]

OpenStudy (aum):

\[ \frac{ \cancel{n!} }{ (n-r)!r! } \times \frac{ (n-r+1)!(r+1)! }{\cancel{n!}} = 1 \\ \frac{ (n-r+1)*(n-r)! *(r+1)*r! }{(n-r)!r!} = 1 \\ \]

OpenStudy (aum):

No. The above is based on the previous result which was wrong.

OpenStudy (anonymous):

wait how did you (n-r)! on top

OpenStudy (aum):

\[ ^nC_r = ^nC_{r+1} = 126 \\ \frac{n!}{(n-r)!*r!} = \frac{n!}{(n-r-1)!*(r+1)!} = 126 \\ \frac{n!}{(n-r)!*r!} \div \frac{n!}{(n-r-1)!*(r+1)!} = 1 \\ \frac{n!}{(n-r)!*r!} \times \frac{(n-r-1)!*(r+1)!}{n!} = 1 \\ \frac{(n-r-1)!*(r+1)!}{(n-r)!*r!} = 1 \\ \frac{(n-r-1)!*(r+1) * r!}{(n-r) * (n-r-1)!*r!} = 1 \\ \frac{r+1}{n-r} = 1 \\ r + 1 = n - r \\ 2r = n - 1 \\ r = \frac 12 (n-1) \]

OpenStudy (aum):

I think Pascal's triangle is the easier method.

OpenStudy (anonymous):

now that i see that, I think the same lol

OpenStudy (anonymous):

Thanks for the help, Pascal's triangle is the way to go for this it seems.

OpenStudy (aum):

You are welcome.